# The non-Hermitian skin effect: A perspective

### v1

@@ Abstract
The non-Hermitian (NH) skin effect is a truly NH feature, which manifests itself as an accumulation of states, known as skin states, on the boundaries of a system. In this perspective, we discuss several aspects of the NH skin effect focusing on the most interesting facets of this phenomenon. Non-normality and non-reciprocity are reviewed as necessary requirements to see the NH skin effect. We further discuss the NH skin effect as a topological effect that can be seen as a manifestation of a truly NH bulk-boundary correspondence, and show how topological boundary states can be distinguished from skin states. As most theoretical work has focused on studying the NH skin effect in the one-dimensional single-particle picture, recent developments of studying this effect in higher dimensions as well as in the many-body case are also highlighted. Lastly, experimental realizations and applications of the NH skin effect are reviewed.
@@ 1 Introduction
Non-Hermitian (NH) topology has established itself as a new exciting research domain over the last decade [1, 2, 3]. Originally motivated from the fact that parity-time (PT) symmetric systems have a completely real eigenspectrum in the symmetry-unbroken regime [4, 5] and the subsequent surge in optical experiments [6, 7, 8, 9, 10], non-Hermiticity is increasingly studied in a wide variety of fields in physics ranging from mechanical [11], electrical [12], acoustic [13] and open quantum systems [14, 15] to strongly correlated phases of matter [16, 17, 18, 19, 20, 21, 22, 23] and non-conservative biological systems [24, 25]. In the field of NH topology, non-Hermiticity is studied through the lens of condensed matter physics uncovering a remarkable enrichment of topological phenomena [1].
At the core of this enrichment is the fact that not only the eigenvectors but also the eigenvalues may have non-trivial topological features. Indeed, the spectrum now lives on the complex plane, such that eigenvalues may wind in a non-trivial manner captured by a spectral invariant [26, 27, 28, 29] as illustrated in Fig. 1. It has been shown that topologically non-trivial spectral features in the eigenvalue spectrum under periodic boundary conditions (PBCs) result in macroscopic accumulation of states on the boundary of a lattice model under open boundary conditions (OBCs) [30, 31, 28, 32, 29, 33] thus establishing a truly non-Hermitian bulk-boundary correspondence (BBC). These exponentially localized states are the so-called NH skin states, and the phenomenon over all is known as the non-Hermitian skin effect, a term coined in Ref. 30, which is the central subject of this perspective.
Here we provide an overview of what we believe are the most interesting aspects of the NH skin effect. Besides a detailed discussion of the aforementioned NH BBC, we discuss how non-normality and non-reciprocity of NH matrices are necessary ingredients to observe the NH skin effect [34]. We continue by explicitly showing how the biorthogonal properties [35] of skin states allow us to distinguish them from topological boundary states. As the NH skin effect has predominantly been studied in the one-dimensional, single-particle case, we also discuss the recent developments in expanding the theory to higher dimensions with a focus on the amoeba formalism [36] as well as to the many-body context [19, 20, 18, 22, 21]. Moving away from pure theory, we continue by providing a short overview of experimental realizations of the NH skin effect as well as exciting applications of it, namely, the funneling of light [9], the topological Ohmmeter [37, 38], and topological amplification [39, 40]. We end the perspective by discussing what we believe are the open questions in the field.
Figure 1: Hatano-Nelson model and its features. (a) Schematic illustration of the model, showing the localization of bulk modes under OBC on the left or right end depending on the relative hopping strengths: For $|J_{L}|<|J_{R}|$, modes localize on the right end, while for $|J_{L}|>|J_{R}|$, they localize on the left end. (b) Under PBCs, the eigenvalues trace an ellipse in the complex plane, and the eigenstates are extended. In contrast, the OBC spectrum drastically deviates from the PBC spectrum exhibiting purely real eigenvalues and localized eigenmodes manifesting the NH skin effect. (c) The real and imaginary parts of the eigenvalues as a function of $k$. The gray and black crosses in (b) and (c) indicate the left and right movers in the dispersion, respectively.
@@ 2 The NH skin effect: The basics
We start this perspective with a paradigmatic one-dimensional (1D) example, and discuss several well-known, yet important general properties of the NH skin effect in 1D systems.
@@ 2.1 The Hatano-Nelson model
with $J_{L}(J_{R})\in\mathbb{R}$ the nearest-neighbor hopping parameter to the left (right), and $c^{\dagger}_{n}$ ($c_{n}$) creating (annihilating) an excitation on site $n$. This model is NH when $J_{L}\neq J_{R}$. As a 1D single-band model, the energy under PBCs, $E(k)$, corresponds to the Bloch Hamiltonian $H(k)$, and reads $E(k)=(J_{L}+J_{R})\cos k+i(J_{L}-J_{R})\sin k$. As seen in Fig. 1(b), $E(k)$ forms an ellipse and winds around the origin in the complex energy plane as a function of $k$ in the (counter)clockwise direction, when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$). As such, it is possible to define a spectral winding number $w$ as [27, 26]
$w=\frac{1}{2\pi i}\int_{-k}^{k}\mathrm{d}k\,\partial_{k}\ln E(k)=\begin{cases}+1,&|J_{L}|>|J_{R}|,\\ -1,&|J_{L}|<|J_{R}|.\end{cases}$ (2)
The winding number can only change when $E(k)=0$ for some $k$, i.e., when $|J_{L}|=|J_{R}|$. Turning to OBCs, it is straightforward to see from the Hamiltonian in Eq. (1), that all eigenstates accumulate on the boundary on the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$) as shown in Fig. 1(a). This phenomenon of a piling up of a macroscopic number of states is known as the NH skin effect.
There is an intuitive way to understand the connection between the sign of the spectral winding number and the boundary to which the NH skin states localize. When the winding number is negative (positive), the eigenvalues wind in the (counter)clockwise direction. As the PBC spectrum of the Hatano-Nelson model is centered around zero energy, half of the spectrum is in the positive imaginary plane. That means that for a negative (positive) winding number, the right (left) movers with group velocity $v_{k}=\mathop{\mathrm{Re}}[\partial_{k}E(k)]>0$ ($<0$) have positive imaginary energy whereas the left (right) movers have negative imaginary energy, as illustrated in Figs. 1(b,c). In the long-time limit, the right (left) movers thus dominate the behavior of the system, and one sees that all NH skin states pile up on the right (left) boundary of the system. A similar argument is also presented in Ref. 42.
@@ 2.2 Localization of skin states
To observe the NH skin effect, one has to study the localization of the right or left eigenstate separately. Indeed, it is shown in Refs. 43, 44, 33 that if the right eigenstates |$\ket{\Psi_{R}}$, which is obtained by solving $H\ket{\Psi_{R}}=E\ket{\Psi_{R}}$, are exponentially localized to one boundary, the left eigenstates |$\ket{\Psi_{L}}$ satisfying $\bra{\Psi_{L}}H=E\bra{\Psi_{L}}$ |, or alternatively, $H^{\dagger}\ket{\Psi_{L}}=E^{*}\ket{\Psi_{L}}$, are exponentially localized to the opposite boundary. As a consequence, the biorthogonal localization of the right and left eigenstate quantified by $\braket{\Psi_{L}}{\Pi_{n}}{\Psi_{R}}$, where $\Pi_{n}=c_{n}^{\dagger}\ket{0}\bra{0}c_{n}$ is the projector onto each site $n$, has weight throughout the system. Thus, the skin states have a bulk-state-like fingerprint in the biorthogonal picture, albeit with potentially negative and/or complex weights.
@@ 2.3 Generalized Brillouin zone
Next, we note there is an intimate link between the NH skin effect and the so-called generalized Brillouin zone (GBZ). The theory of the GBZ was first developed by Yao and Wang in Ref. 30, and later expanded in Refs. 45, 44, in order to restore the breakdown of the conventional bulk-boundary correspondence (cBBC) as observed in models featuring the NH skin effect, which will be discussed in more detail below.
The GBZ $\mathcal{C}$ can be constructed by solving the characteristic equation of the non-Bloch Hamiltonian $H(\beta)$, which is the analytic continuation of the Bloch Hamiltonian. This closed curve $\mathcal{C}$ contains crucial information about the system under OBCs: The OBC spectrum is given by $H(\beta)$ with $\beta\in\mathcal{C}$, and one can determine spatial decay rates for the NH skin modes, which are related to $\beta$, in the continuum limit. Indeed, for any skin mode, one finds its associated $\beta$, and if $|\beta|<1$ $(>1)$, or in other words if $\mathcal{C}$ is inside (outside) of the unit circle [45], the skin state localize to the left (right) boundary. For $|\beta|=1$, the state is delocalized as the Bloch states, and the corresponding point is called a Bloch point [45].
@@ 2.4 Exceptional points
Interestingly, the NH skin effect is accompanied by the appearance of exceptional points (EPs) with an order scaling with system size under OBCs [46, 47]. EPs are degeneracies, which abundantly appear in NH matrices [48, 49], at which not only the eigenvalues but also the eigenvectors coalesce, i.e., the geometric multiplicity is smaller than the algebraic multiplicity [50, 51]. At EPs, the NH matrix features Jordan blocks, whose dimension corresponds to the order of the EP. Close to an EP, eigenvectors start to overlap until they finally coalesce when reaching the EP. We indeed see such a high-order EP for the Hatano-Nelson model, where in the extreme limit $J_{R}\,(J_{L})=0$, all states coalesce onto the last left (right) site .
@@ 2.5 Sensitivity to boundary conditions
As is clearly visible in the spectrum of the Hatano-Nelson model in Fig. 1(b), the spectrum strongly depends on the boundary conditions. In Ref. 52, it is shown that by introducing coupling terms between the ends of 1D chains, one finds that the system crosses several EPs upon tuning tune between OBCs and PBCs [53, 54, 55], highlighting a topological distinction between the two cases. Refs. 53, 54 show that the cross-over from OBCs to PBCs, i.e., from having skin states to Bloch states, happens for exponentially small couplings. While it thus may seem that the NH skin effect is not a robust feature, it is shown in Ref. 54 that physically relevant locality constraints exponentially suppress this kind of coupling. Also, further modifications to the boundary conditions result in new localization phenomena and transitions between different behaviors [56].
@@ 2.6 Beyond Hamiltonians
While usually phrased in terms of Hamiltonians, the skin effect is a property of NH matrices. Thus, any physical system which can be represented by a NH matrix might exhibit the skin effect. Prominent examples include the conductance matrix, which is the appropriate object to consider when handling topoelectric circuits [12], and the damping matrix in open quantum systems [14, 15].
@@ 3 Non-normality and non-reciprocity as requirement of the NH skin effect
Not all NH matrices feature a NH skin effect, and a necessary criterion needs to be satisfied, namely, the NH matrix needs to be non-normal [1], i.e., $[H,H^{\dagger}]\neq 0$. By writing $H=H_{\text{H}}+iH_{\text{A}}$, with $H_{\text{H}}=H_{\text{H}}^{\dagger}$ the Hermitian part and $iH_{\text{A}}=-(iH_{\text{A}})^{\dagger}$ the anti-Hermitian part of $H$, non-normality is equivalent to $[H_{\text{H}},H_{\text{A}}]\neq 0$. If $H_{\text{H}}$ and $H_{\text{A}}$ would commute, $H$ would share an orthogonal eigenbasis with $H_{\text{H}}$ and $H_{\text{A}}$, which prohibits the existence of skin states.
In Ref. 34, it is shown that non-normality is not a sufficient condition. Additionally, the system needs to be non-reciprocal, that is, if its susceptibility matrix, or Green’s function, $\chi(\omega)=-i(\omega\mathbb{I}-H)^{-1}$ satisfies $|\chi(\omega)|\neq|\chi(\omega)|^{T}$, which is equivalent to the scattering matrix $S$ satisfying $|S|\neq|S|^{T}$, where the modulus of a matrix means the absolute value of each matrix element [57, 58]. Physically, this means that the response of the system is not invariant under exchanging input and output. Importantly, this presented notion of non-reciprocity should not be confused with non-reciprocal hoppings, which in the NH context refers to asymmetric hoppings and in the Hermitian context is sometimes used to indicate the presence of non-zero phases on the hoppings due to pseudo-magnetic fields.
Figure 2: Topological end state and skin state localization. Normalized right wavefunction of (a) topological boundary state in pink and (b) skin states in blues. The end state is localized on a single sublattice whereas the skin modes are distributed across both sublattices. (c) The biorthogonal product of the right and left states shows that the end state remains localized while the skin states are delocalized.
@@ 4 Spectral features, topological protection and the NH bulk-boundary correspondence
If a NH matrix is both non-normal and non-reciprocal, its PBC spectrum displays a so-called point gap [34]. A point gap is defined as there being a base point $E_{B}$, which is not crossed by the complex-energy bands, and crossing this point defines a gap closing transition [27]. Indeed, we see in the example of the Hatano-Nelson model that the PBC spectrum has a point gap, cf. Fig. 1(b), and the gap closes as discussed at $|J_{L}|=|J_{R}|$. It has been shown that for the NH skin effect to occur in 1D systems, the PBC spectrum must feature such point gaps [28, 32, 59, 33]. In extension, this means that it is possible to find a non-zero spectral topological invariant for these systems. For example, the spectral winding number in Eq. (2) for arbitrary periodic 1D systems described by the Bloch Hamiltonian $H(k)$, can be generalized to the following form [27]
@@
We note that if one is interested in the state localization of a specific skin state, one has to select its energy as reference energy $E_{B}$.
The skin effect as a topological effect is robust against weak perturbations of the system. For example, in the Hatano-Nelson model, cf. Eq. (1), onsite disorder has been studied in Refs. 41, 27, whereas additional perturbations to the hoppings have been addressed in Ref. 60, 61, revealing a stability of the skin states to these disturbances.

### v2

@@ Abstract
The non-Hermitian (NH) skin effect is a truly NH feature, which manifests itself as an accumulation of states, known as skin states, on the boundaries of a system. In this perspective, we discuss several aspects of the NH skin effect focusing on the most interesting facets of this phenomenon. Beyond reviewing necessary requirements to see the NH skin effect, we discuss the NH skin effect as a topological effect that can be seen as a manifestation of a truly NH bulk-boundary correspondence, stemming from the spectral topology, and show how skin states can be distinguished from topological boundary states. As most theoretical work has focused on studying the NH skin effect in one-dimensional non-interacting systems, recent developments of studying this effect in higher dimensions as well as in many-body systems are highlighted. Lastly, experimental signatures and applications are discussed, and an outlook is provided.
@@ 1 Introduction
Non-Hermitian (NH) topology has established itself as a new exciting research domain over the last decade [1, 2, 3]. Originally motivated from the fact that parity-time symmetric systems may have a completely real eigenspectrum [4, 5], and the subsequent surge in optical experiments [6, 7, 8, 9, 10], non-Hermiticity is increasingly studied in a wide variety of fields in physics ranging from mechanical [11], electrical [12], acoustic [13] and open quantum systems [14, 15] to quantum scattering [16], strongly correlated phases of matter [17, 18, 19, 20, 21, 22, 23, 24] and non-conservative biological systems [25, 26]. In the field of NH topology, non-Hermiticity is studied through the lens of condensed matter physics uncovering a remarkable enrichment of topological phenomena [1].
At the core of this enrichment is the fact that not only the eigenvectors, but also the eigenvalues may have non-trivial topological features. Indeed, the spectrum now lives on the complex plane, such that eigenvalues may wind in a non-trivial manner captured by a spectral invariant [27, 28, 29, 30, 31] as illustrated in Fig. 1. It has been shown that topologically non-trivial spectral features under periodic boundary conditions (PBCs) result in a macroscopic accumulation of states on the boundary of a lattice model under open boundary conditions (OBCs) [32, 33, 30, 34, 31, 35] establishing a truly non-Hermitian bulk-boundary correspondence. These exponentially localized states are the so-called NH skin states, and the phenomenon overall is known as the non-Hermitian skin effect, a term coined in Ref. 32, which is the central subject of this perspective.
Figure 1: Hatano-Nelson model and its features. (a) Schematic illustration of the model showing the localization of bulk modes under OBCs on the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$). (b) Under PBCs, the eigenvalues trace an ellipse in the complex plane (green), and the eigenstates are extended. In contrast, the OBC spectrum (purple) drastically deviates from the PBC spectrum exhibiting purely real eigenvalues and localized eigenmodes manifesting the NH skin effect. (c) The real (solid) and imaginary (dashed) parts of the PBC energies as a function of $k$. The gray and black crosses in (b) and (c) indicate the left and right movers in the dispersion, respectively. $J_{L}=3/2$ and $J_{R}=1/2$ in (b) and (c).
@@ 2 The NH skin effect: The basics
We start this perspective with a paradigmatic one-dimensional (1D) example, as to discuss several general properties we believe are well-known, yet important to understand the NH skin effect in 1D systems.
@@ 2.1 The Hatano-Nelson model
with $J_{L}\,(J_{R})\in\mathbb{R}$ the nearest-neighbor (NN) hopping parameter to the left (right), and $c^{\dagger}_{n}$ ($c_{n}$) creating (annihilating) an excitation on site $n$. This model is NH when $J_{L}\neq J_{R}$. As a 1D single-band model, the energy under PBCs $E(k)$ corresponds to the Bloch Hamiltonian $H(k)$, and reads $E(k)=(J_{L}+J_{R})\cos k+i(J_{L}-J_{R})\sin k$. As seen in Fig. 1(b), $E(k)$ forms an ellipse and winds around the origin in the complex energy plane as a function of $k$ in the (counter)clockwise direction, when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$). As such, it is possible to define a spectral winding number $w$ as [27, 29, 28]
$w=\frac{1}{2\pi i}\int_{-\pi}^{\pi}\mathrm{d}k\,\partial_{k}\ln E(k)=\begin{cases}+1,&|J_{L}|>|J_{R}|,\\ -1,&|J_{L}|<|J_{R}|.\end{cases}$ (2)
The winding number can only change when $E(k)=0$ for some $k$, i.e., when $|J_{L}|=|J_{R}|$. Turning to OBCs, it is straightforward to see from the Hamiltonian in Eq. (1), that all eigenstates satisfying $H\ket{\Psi_{R}}=E\ket{\Psi_{R}}$ are exponentially localized on the boundary on the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$) as shown in Fig. 1(a). This phenomenon of a piling up of a macroscopic number of states on the boundary is the NH skin effect.
We find there is an intuitive and enlightening way to understand the connection between the sign of the spectral winding number and the boundary to which the skin states localize. When the winding number is negative (positive), the eigenvalues wind in the (counter)clockwise direction. As the PBC spectrum of the Hatano-Nelson model is centered around zero energy, half of the spectrum is in the positive imaginary plane. That means that for a positive winding number, the right (left) movers with group velocity $v_{k}=\mathop{\mathrm{Re}}[\partial_{k}E(k)]>0$ ($<0$) have negative (positive) imaginary energy, as illustrated in Figs. 1(b,c). In the long-time limit, the left movers thus dominate the behavior of the system, and one sees that all NH skin states pile up on the left boundary of the system. The reverse holds for negative winding number. A similar argument is presented in Ref. 37.
@@ 2.2 Localization of skin states
When considering eigenstates of a NH system, one should take care of the appropriate biorthogonalization [38]: A NH system does not only have right eigenstates |$\ket{\Psi_{R}}$, but also distinct left eigenvectors satisfying $\bra{\Psi_{L}}H=E\bra{\Psi_{L}}$ |, equivalent to $H^{\dagger}\ket{\Psi_{L}}=E^{*}\ket{\Psi_{L}}$. By studying the left and right eigenstates separately one can observe the skin effect. Indeed, if |$\ket{\Psi_{R}}$ is localized on one boundary, the corresponding $\bra{\Psi_{L}}$ | is localized to the opposite boundary [39, 40, 35]. As a consequence, the biorthogonal localization of the right and left eigenstate quantified by $\braket{\Psi_{L}}{\Pi_{n}}{\Psi_{R}}$, where $\Pi_{n}=c_{n}^{\dagger}\ket{0}\bra{0}c_{n}$ is the projector onto each site $n$, has weight throughout the system. Thus, skin states have a bulk-state-like fingerprint in the biorthogonal picture, albeit with potentially negative or complex weights.
@@ 2.3 Generalized Brillouin zone
Next, we point out the intimate link between the NH skin effect and the so-called generalized Brillouin zone (GBZ). The theory of the GBZ was first developed by Yao and Wang in Ref. 32, and later expanded in Refs. 41 and 40. The GBZ is a closed curve $\mathcal{C}$ solving the characteristic equation of the non-Bloch Hamiltonian $H(\beta)$, which is the analytic continuation of the Bloch Hamiltonian. The GBZ contains crucial information about the system under OBCs: The OBC spectrum is given by $H(\beta)$ with $\beta\in\mathcal{C}$, and one can determine spatial decay rates for the NH skin modes, which are related to $\beta$, in the continuum limit. Indeed, for any skin mode, one finds its associated $\beta$, and if $|\beta|<1$ $(>1)$, or in other words if $\mathcal{C}$ is inside (outside) of the unit circle [41], the skin state localize to the left (right) boundary. For $|\beta|=1$, the state is delocalized as the Bloch states, and the corresponding point is called a Bloch point [41].
@@ 2.4 Exceptional points
Interestingly, the NH skin effect is accompanied by the appearance of exceptional points (EPs) with an order scaling with system size under OBCs [42, 43]. EPs are degeneracies, which abundantly appear in NH matrices [44, 45], at which not only the eigenvalues but also the eigenvectors coalesce, i.e., the geometric multiplicity is smaller than the algebraic multiplicity [46, 47]. At EPs, the NH matrix features Jordan blocks, whose dimension corresponds to the order of the EP. Close to an EP, eigenvectors start to overlap until they finally coalesce when reaching the EP. We indeed see such a high-order EP for the Hatano-Nelson model, where in the extreme limit $J_{R}\,(J_{L})=0$, all states coalesce onto a single one with weight only on the leftmost (rightmost) site.
@@ 2.5 Sensitivity to boundary conditions
Many questions often arise in relation to the thermodynamic limit in NH systems. Here we highlight how subtle this matter is by studying the sensitivity to boundary conditions in the Hatano-Nelson model. As is clearly visible in the spectrum in Fig. 1(b), it strongly depends on the boundary conditions. In Ref. 48, it is shown that by introducing coupling terms between the ends of 1D chains, one finds that the system crosses several EPs upon tuning tune between OBCs and PBCs [49, 50, 51], emphasizing a topological distinction between the two cases. Refs. 49 and 50 show that the cross-over from OBCs to PBCs, i.e., from having skin states to Bloch states, happens for exponentially small couplings. While it thus may seem that the NH skin effect is not a robust feature, it is shown in Ref. 50 that physically relevant locality constraints exponentially suppress this kind of coupling. Also, further modifications to the boundary conditions result in new localization phenomena and transitions between different behaviors [52].
@@ 2.6 Beyond Hamiltonians
While usually phrased in terms of Hamiltonians, we note that the skin effect is a property of NH matrices or operators. Thus, any physical system represented by such may exhibit the skin effect. Examples include the conductance matrix, which is the appropriate object to consider when handling topoelectric circuits [12], the damping matrix in open quantum systems [14, 15], and the scattering matrix [16, 53].
@@ 3 Non-normality and non-reciprocity as requirement of the NH skin effect
We want to point out that not all NH matrices feature the NH skin effect, and a necessary criterion needs to be satisfied, namely, the NH matrix needs to be non-normal [1], i.e., $[H,H^{\dagger}]\neq 0$. By writing $H=H_{\text{H}}+iH_{\text{A}}$, with $H_{\text{H}}=H_{\text{H}}^{\dagger}$ the Hermitian part and $iH_{\text{A}}=-(iH_{\text{A}})^{\dagger}$ the anti-Hermitian part of $H$, non-normality is equivalent to $[H_{\text{H}},H_{\text{A}}]\neq 0$. If $H_{\text{H}}$ and $H_{\text{A}}$ would commute, $H$ would share an orthogonal eigenbasis with $H_{\text{H}}$ and $H_{\text{A}}$, which prohibits the existence of skin states.
Additionally, a system needs to be non-reciprocal to show the NH skin effect [53]. This can either be checked by analyzing if the systems Green’s function, $\chi(\omega)=-i(\omega\mathbb{I}-H)^{-1}$, satisfies $|\chi(\omega)|\neq|\chi(\omega)|^{T}$, where the modulus of a matrix means the absolute value of each matrix element, or by analyzing the systems’ singular value decomposition (SVD) [53]. In the scattering context discussed in Ref. 53, the requirement of non-reciprocity is equivalent to the requirement that the systems scattering matrix $S$ satisfies $|S|\neq|S|^{T}$. Physically, this means that the response of the system is not invariant under exchanging input and output [54, 55]. While non-normality and non-reciprocity are equivalent for single-band NN models, e.g., the aforementioned Hatano-Nelson model, this is in general not true for single-band systems including longer-range hoppings and multi-band systems [53].
We want to emphasize that the presented definition of non-reciprocity is different from the notion of non-reciprocal hoppings, i.e, hoppings whose magnitude depends on the hopping direction, which we refer to as asymmetric hoppings. It is namely possible to observe the skin effect without asymmetric hoppings: While the frequently studied anisotropic Su-Schrieffer-Heeger (SSH) model in Eq. (4) below features asymmetric hoppings, it is unitarily equivalent to the NH Creutz ladder [1] introduced by Lee as a first example for the breakdown of the conventional bulk-boundary correspondence (cBBC) [56, 1]. The NH Creutz ladder does not exhibit asymmetric hoppings, but instead features on-site gain and loss as well as phases on the hoppings. Nevertheless, this model is non-reciprocal in the aforementioned sense and shows the NH skin effect.
Figure 2: Topological end state and skin state localization. Normalized right eigenstate of (a) topological end state in pink and (b) skin states in blues for the model in Eq. (4). The end state is localized on a single sublattice whereas the skin modes are distributed across both sublattices. (c) The biorthogonal product of the right and left states shows that the end state remains localized while the skin states are delocalized. We fix $t_{1}=3/2,t_{2}=2,$ $\gamma=0.14$, and $N=41$.
@@ 4 Spectral features, topological protection and the NH bulk-boundary correspondence
If a NH matrix is both non-normal and non-reciprocal, its PBC spectrum displays a so-called point gap [53]. A point gap is defined as there being a base point $E_{B}$, which is not crossed by the complex-energy bands, and crossing this point defines a gap closing transition [29]. Indeed, we see in the example of the Hatano-Nelson model that the PBC spectrum has a point gap, cf. Fig. 1(b), and the gap closes as discussed at $|J_{L}|=|J_{R}|$. It has been shown that for the NH skin effect to occur in 1D systems, the PBC spectrum must feature such point gaps [30, 34, 57, 35]. In extension, this means that it is possible to find a non-zero spectral topological invariant for these systems. For example, the spectral winding number in Eq. (2) for arbitrary periodic 1D systems described by the Bloch Hamiltonian $H(k)$, can be generalized to the following form [29]:
@@
We note that if one is interested in the state localization of a specific skin state, one has to select its energy as reference energy $E_{B}$. While it is quite intuitive that all eigenstates are skewed towards the right (left) when $|J_{L}|<|J_{R}|$ ($>|J_{R}|$) in the Hatano-Nelson model, the spectral winding number is essential to understand the competition between different hopping strengths and ranges in more general models.
The skin effect as a topological effect is robust against weak perturbations of the system. For example, the Hatano-Nelson model, cf. Eq. (1), is robust against weak onsite disorder [36, 29] and additional perturbations to the hoppings [58, 59], revealing a stability of the skin states to these disturbances.